OpenStudy (anonymous):

Okay, so I know that simple interest is an arithmetic sequence, T(n) = a + (n-1)d, but cant for the life of me, start at T(0). I am using the example of £100 invested at APR of 5%, and trying to calculate T(0), T(1), T(2) etc.. maybe I can only start at T(1), can anyone answer?

OpenStudy (perl):

the formula T(n) = a +(n-1)*d implicitly starts at n=1

OpenStudy (perl):

if you want to start at n = 0, you can just use the formula T(n) = a + n*d

OpenStudy (perl):

for simple interest: Balance = P ( 1 + r*n ) where r = annual interest, n = number of years, P = principal using T(n) for balance, we have T(n) = 100 ( 1 + .05*n) = 100 * 1 + 100*.05*n = 100 + 5*n so start at n= 0 T(n) = 100 + 5n

OpenStudy (anonymous):

Is the form T(n) = a + n*d the closed form for t(0)?

OpenStudy (anonymous):

Yes, I started with a= 100 euro, and then obviously T1 = a + interest but my problem was: to be an arithmetic sequence it has to fit the formula: T(n) = a + (n-1)d which obviously this doesnt!